OpenStudy (anonymous):

HELP ME PLEASE! Using your function, explain how to solve for f(3). Show your work and explain each step using complete sentences.

OpenStudy (anonymous):

what is your function?

OpenStudy (anonymous):

I'm not sure, I never really made one, but the equation i've been solving is: 2x + 4 = –3x + 14 I suppose this could be a function

OpenStudy (anonymous):

no it cannot a function is not an equation, although it can be defined by one a function needs to look like \(f(x)=\text{expression}\)

OpenStudy (anonymous):

for example you could have \[f(x)=2x+4\] or \[f(x)=x^2+1\] or \[f(x)=\frac{1}{x}\] or whatever

OpenStudy (anonymous):

Alright, so how would I work f(3) into this?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@dmezzullo